# Universal neural network potentials as descriptors: Towards scalable chemical property prediction using quantum and classical computers

### v1

@@ Acknowledgements.
We thank Nobuki Inoue and Tuan Minh Do for fruitful discussions. This project was supported by funding from the MEXT Quantum Leap Flagship Program (MEXTQLEAP) through Grant No. JPMXS0120319794, and the JST COI-NEXT Program through Grant No. JPMJPF2014. The completion of this research was partially facilitated by the JSPS Grants-in-Aid for Scientific Research (KAKENHI), specifically Grant Nos. JP23H03819 and JP21K18933. We thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was also achieved through the use of SQUID at the Cybermedia Center, Osaka University.References
- Reymond et al. [2010] J.-L. Reymond, R. van Deursen, L. C. Blum, and L. Ruddigkeit, Chemical space as a source for new drugs, MedChemComm 1, 30 (2010).
@@
- Ruddigkeit et al. [2012] L. Ruddigkeit, R. van Deursen, L. C. Blum, and J.-L. Reymond, Enumeration of 166 billion organic small molecules in the chemical universe database GDB-17, Journal of Chemical Information and Modeling 52, 2864 (2012).
@@
- Gómez-Bombarelli et al. [2016] R. Gómez-Bombarelli, J. Aguilera-Iparraguirre, T. D. Hirzel, D. Duvenaud, D. Maclaurin, M. A. Blood-Forsythe, H. S. Chae, M. Einzinger, D.-G. Ha, T. Wu, G. Markopoulos, S. Jeon, H. Kang, H. Miyazaki, M. Numata, S. Kim, W. Huang, S. I. Hong, M. Baldo, R. P. Adams, and A. Aspuru-Guzik, Design of efficient molecular organic light-emitting diodes by a high-throughput virtual screening and experimental approach, Nature Materials 15, 1120 (2016).
@@
- Curtarolo et al. [2013] S. Curtarolo, G. L. W. Hart, M. B. Nardelli, N. Mingo, S. Sanvito, and O. Levy, The high-throughput highway to computational materials design, Nature Materials 12, 191 (2013).
- Nie et al. [2023] W. Nie, Q. Wan, J. Sun, M. Chen, M. Gao, and S. Chen, Ultra-high-throughput mapping of the chemical space of asymmetric catalysis enables accelerated reaction discovery, Nature Communications 14, 10.1038/s41467-023-42446-5 (2023).
- Oganov et al. [2019] A. R. Oganov, C. J. Pickard, Q. Zhu, and R. J. Needs, Structure prediction drives materials discovery, Nature Reviews Materials 4, 331 (2019).
- Pierens [2014] G. K. Pierens, 1h and 13c nmr scaling factors for the calculation of chemical shifts in commonly used solvents using density functional theory, Journal of computational chemistry 35, 1388 (2014).
@@
- Hartman et al. [2016] J. D. Hartman, R. A. Kudla, G. M. Day, L. J. Mueller, and G. J. O. Beran, Benchmark fragment-based 1h, 13c, 15n and 17o chemical shift predictions in molecular crystals, Physical Chemistry Chemical Physics 18, 21686 (2016).
- Büning and Grimme [2023] J. B. K. Büning and S. Grimme, Computation of CCSD(t)-quality NMR chemical shifts via machine learning from DFT, Journal of Chemical Theory and Computation 19, 3601 (2023).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
- Lodewyk et al. [2011] M. W. Lodewyk, M. R. Siebert, and D. J. Tantillo, Computational prediction of 1h and 13c chemical shifts: A useful tool for natural product, mechanistic, and synthetic organic chemistry, Chemical Reviews 112, 1839 (2011).
- Lauro et al. [2020] G. Lauro, P. Das, R. Riccio, D. S. Reddy, and G. Bifulco, Dft/nmr approach for the configuration assignment of groups of stereoisomers by the combination and comparison of experimental and predicted sets of data, The Journal of Organic Chemistry 85, 3297–3306 (2020).
@@
- Hansen et al. [2015] K. Hansen, F. Biegler, R. Ramakrishnan, W. Pronobis, O. A. Von Lilienfeld, K.-R. Muller, and A. Tkatchenko, Machine learning predictions of molecular properties: Accurate many-body potentials and nonlocality in chemical space, The journal of physical chemistry letters 6, 2326 (2015).
- Deringer et al. [2021] V. L. Deringer, A. P. Bartók, N. Bernstein, D. M. Wilkins, M. Ceriotti, and G. Csányi, Gaussian process regression for materials and molecules, Chemical Reviews 121, 10073 (2021).
@@
- Faber et al. [2017] F. A. Faber, L. Hutchison, B. Huang, J. Gilmer, S. S. Schoenholz, G. E. Dahl, O. Vinyals, S. Kearnes, P. F. Riley, and O. A. von Lilienfeld, Prediction errors of molecular machine learning models lower than hybrid dft error, Journal of Chemical Theory and Computation 13, 5255 (2017).
@@
- Sajjan et al. [2022] M. Sajjan, J. Li, R. Selvarajan, S. H. Sureshbabu, S. S. Kale, R. Gupta, V. Singh, and S. Kais, Quantum machine learning for chemistry and physics, Chemical Society Reviews 51, 6475 (2022).
@@
- Keith et al. [2021] J. A. Keith, V. Vassilev-Galindo, B. Cheng, S. Chmiela, M. Gastegger, K.-R. Müller, and A. Tkatchenko, Combining machine learning and computational chemistry for predictive insights into chemical systems, Chemical Reviews 121, 9816 (2021).
- Wan et al. [2023] K. Wan, J. He, and X. Shi, Construction of high accuracy machine learning interatomic potential for surface/interface of nanomaterials—a review, Advanced Materials 10.1002/adma.202305758 (2023).
@@
- Wu et al. [2018] Z. Wu, B. Ramsundar, E. N. Feinberg, J. Gomes, C. Geniesse, A. S. Pappu, K. Leswing, and V. Pande, MoleculeNet: a benchmark for molecular machine learning, Chemical Science 9, 513 (2018).
@@
- Reiser et al. [2022] P. Reiser, M. Neubert, A. Eberhard, L. Torresi, C. Zhou, C. Shao, H. Metni, C. van Hoesel, H. Schopmans, T. Sommer, and P. Friederich, Graph neural networks for materials science and chemistry, Communications Materials 3, 10.1038/s43246-022-00315-6 (2022).
- Bartók et al. [2017] A. P. Bartók, S. De, C. Poelking, N. Bernstein, J. R. Kermode, G. Csányi, and M. Ceriotti, Machine learning unifies the modeling of materials and molecules, Science Advances 3, 10.1126/sciadv.1701816 (2017).
- Kocer et al. [2022] E. Kocer, T. W. Ko, and J. Behler, Neural network potentials: A concise overview of methods, Annual Review of Physical Chemistry 73, 163–186 (2022).
- Langer et al. [2022] M. F. Langer, A. Goeßmann, and M. Rupp, Representations of molecules and materials for interpolation of quantum-mechanical simulations via machine learning, npj Computational Materials 8, 41 (2022).
- Liu et al. [2021] Z. Liu, L. Lin, Q. Jia, Z. Cheng, Y. Jiang, Y. Guo, and J. Ma, Transferable multilevel attention neural network for accurate prediction of quantum chemistry properties via multitask learning, Journal of Chemical Information and Modeling 61, 1066–1082 (2021).
- Merchant et al. [2023a] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 624, 80–85 (2023a).
- Gao et al. [2020] P. Gao, J. Zhang, Q. Peng, J. Zhang, and V.-A. Glezakou, General protocol for the accurate prediction of molecular 13c/1h nmr chemical shifts via machine learning augmented dft, Journal of Chemical Information and Modeling 60, 3746 (2020).
- Macalino et al. [2015] S. J. Y. Macalino, V. Gosu, S. Hong, and S. Choi, Role of computer-aided drug design in modern drug discovery, Archives of Pharmacal Research 38, 1686 (2015).
@@
- Himanen et al. [2020] L. Himanen, M. O. Jäger, E. V. Morooka, F. Federici Canova, Y. S. Ranawat, D. Z. Gao, P. Rinke, and A. S. Foster, Dscribe: Library of descriptors for machine learning in materials science, Computer Physics Communications 247, 106949 (2020).
- Gupta et al. [2021] A. Gupta, S. Chakraborty, and R. Ramakrishnan, Revving up 13c NMR shielding predictions across chemical space: benchmarks for atoms-in-molecules kernel machine learning with new data for 134 kilo molecules, Machine Learning: Science and Technology 2, 035010 (2021).
@@
- Gerrard et al. [2020] W. Gerrard, L. A. Bratholm, M. J. Packer, A. J. Mulholland, D. R. Glowacki, and C. P. Butts, IMPRESSION – prediction of NMR parameters for 3-dimensional chemical structures using machine learning with near quantum chemical accuracy, Chemical Science 11, 508 (2020).
- Willatt et al. [2019] M. J. Willatt, F. Musil, and M. Ceriotti, Atom-density representations for machine learning, The Journal of Chemical Physics 150, 10.1063/1.5090481 (2019).
- Bartók et al. [2013] A. P. Bartók, R. Kondor, and G. Csányi, On representing chemical environments, Physical Review B 87, 10.1103/physrevb.87.184115 (2013).
- Faber et al. [2018] F. A. Faber, A. S. Christensen, B. Huang, and O. A. von Lilienfeld, Alchemical and structural distribution based representation for universal quantum machine learning, The Journal of Chemical Physics 148, 10.1063/1.5020710 (2018).
@@
- Christensen et al. [2020] A. S. Christensen, L. A. Bratholm, F. A. Faber, and O. A. von Lilienfeld, FCHL revisited: Faster and more accurate quantum machine learning, The Journal of Chemical Physics 152, 10.1063/1.5126701 (2020).
@@
- Kabylda et al. [2023] A. Kabylda, V. Vassilev-Galindo, S. Chmiela, I. Poltavsky, and A. Tkatchenko, Efficient interatomic descriptors for accurate machine learning force fields of extended molecules, Nature Communications 14, 10.1038/s41467-023-39214-w (2023).
- Musil et al. [2021] F. Musil, A. Grisafi, A. P. Bartók, C. Ortner, G. Csányi, and M. Ceriotti, Physics-inspired structural representations for molecules and materials, Chemical Reviews 121, 9759 (2021).
- Bartók et al. [2010] A. P. Bartók, M. C. Payne, R. Kondor, and G. Csányi, Gaussian approximation potentials: The accuracy of quantum mechanics, without the electrons, Physical Review Letters 104, 10.1103/physrevlett.104.136403 (2010).
@@
- Parsaeifard et al. [2021] B. Parsaeifard, D. Sankar De, A. S. Christensen, F. A. Faber, E. Kocer, S. De, J. Behler, O. Anatole von Lilienfeld, and S. Goedecker, An assessment of the structural resolution of various fingerprints commonly used in machine learning, Machine Learning: Science and Technology 2, 015018 (2021).
- Khan et al. [2023] D. Khan, S. Heinen, and O. A. von Lilienfeld, Kernel based quantum machine learning at record rate: Many-body distribution functionals as compact representations, The Journal of Chemical Physics 159, 10.1063/5.0152215 (2023).
- Rupp et al. [2015] M. Rupp, R. Ramakrishnan, and O. A. von Lilienfeld, Machine learning for quantum mechanical properties of atoms in molecules, The Journal of Physical Chemistry Letters 6, 3309 (2015).
@@
- Cordova et al. [2022] M. Cordova, E. A. Engel, A. Stefaniuk, F. Paruzzo, A. Hofstetter, M. Ceriotti, and L. Emsley, A machine learning model of chemical shifts for chemically and structurally diverse molecular solids, The Journal of Physical Chemistry C 126, 16710 (2022).
@@
- Kohlhoff et al. [2009] K. J. Kohlhoff, P. Robustelli, A. Cavalli, X. Salvatella, and M. Vendruscolo, Fast and accurate predictions of protein NMR chemical shifts from interatomic distances, Journal of the American Chemical Society 131, 13894 (2009).
- Tsitsvero et al. [2023] M. Tsitsvero, J. Pirillo, Y. Hijikata, and T. Komatsuzaki, Nmr spectrum prediction for dynamic molecules by machine learning: A case study of trefoil knot molecule, The Journal of Chemical Physics 158, 10.1063/5.0147398 (2023).
- Paruzzo et al. [2018] F. M. Paruzzo, A. Hofstetter, F. Musil, S. De, M. Ceriotti, and L. Emsley, Chemical shifts in molecular solids by machine learning, Nature Communications 9, 10.1038/s41467-018-06972-x (2018).
- Fung et al. [2021] V. Fung, J. Zhang, E. Juarez, and B. G. Sumpter, Benchmarking graph neural networks for materials chemistry, npj Computational Materials 7, 10.1038/s41524-021-00554-0 (2021).
- Guan et al. [2021] Y. Guan, S. V. S. Sowndarya, L. C. Gallegos, P. C. S. John, and R. S. Paton, Real-time prediction of 1h and 13c chemical shifts with DFT accuracy using a 3d graph neural network, Chemical Science 12, 12012 (2021).
@@
- Liu et al. [2019] S. Liu, J. Li, K. C. Bennett, B. Ganoe, T. Stauch, M. Head-Gordon, A. Hexemer, D. Ushizima, and T. Head-Gordon, Multiresolution 3d-DenseNet for chemical shift prediction in NMR crystallography, The Journal of Physical Chemistry Letters 10, 4558 (2019).
- Han and Choi [2021] H. Han and S. Choi, Transfer learning from simulation to experimental data: NMR chemical shift predictions, The Journal of Physical Chemistry Letters 12, 3662 (2021).
- Kwon et al. [2020] Y. Kwon, D. Lee, Y.-S. Choi, M. Kang, and S. Kang, Neural message passing for NMR chemical shift prediction, Journal of Chemical Information and Modeling 60, 2024 (2020).
- Jonas and Kuhn [2019] E. Jonas and S. Kuhn, Rapid prediction of NMR spectral properties with quantified uncertainty, Journal of Cheminformatics 11, 10.1186/s13321-019-0374-3 (2019).
@@
- Schütt et al. [2017] K. T. Schütt, P.-J. Kindermans, H. E. Sauceda, S. Chmiela, A. Tkatchenko, and K.-R. Müller, Schnet: A continuous-filter convolutional neural network for modeling quantum interactions (2017), arXiv:1706.08566 [stat.ML] .
- Chen et al. [2019] C. Chen, W. Ye, Y. Zuo, C. Zheng, and S. P. Ong, Graph networks as a universal machine learning framework for molecules and crystals, Chemistry of Materials 31, 3564 (2019).
- Han et al. [2022] J. Han, H. Kang, S. Kang, Y. Kwon, D. Lee, and Y.-S. Choi, Scalable graph neural network for NMR chemical shift prediction, Physical Chemistry Chemical Physics 24, 26870 (2022).
- Xu et al. [2019] K. Xu, W. Hu, J. Leskovec, and S. Jegelka, How powerful are graph neural networks? (2019), arXiv:1810.00826 [cs.LG] .
@@
- Stärk et al. [2022] H. Stärk, D. Beaini, G. Corso, P. Tossou, C. Dallago, S. Günnemann, and P. Lió, 3D infomax improves GNNs for molecular property prediction, in Proceedings of the 39th International Conference on Machine Learning, Proceedings of Machine Learning Research, Vol. 162, edited by K. Chaudhuri, S. Jegelka, L. Song, C. Szepesvari, G. Niu, and S. Sabato (PMLR, 2022) pp. 20479–20502.
- Jiang et al. [2021] D. Jiang, Z. Wu, C.-Y. Hsieh, G. Chen, B. Liao, Z. Wang, C. Shen, D. Cao, J. Wu, and T. Hou, Could graph neural networks learn better molecular representation for drug discovery? a comparison study of descriptor-based and graph-based models, Journal of Cheminformatics 13, 10.1186/s13321-020-00479-8 (2021).
- Kang et al. [2020] S. Kang, Y. Kwon, D. Lee, and Y.-S. Choi, Predictive modeling of nmr chemical shifts without using atomic-level annotations, Journal of Chemical Information and Modeling 60, 3765–3769 (2020).
@@
- Musaelian et al. [2023] A. Musaelian, S. Batzner, A. Johansson, L. Sun, C. J. Owen, M. Kornbluth, and B. Kozinsky, Learning local equivariant representations for large-scale atomistic dynamics, Nature Communications 14, 579 (2023).
- Merchant et al. [2023b] A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon, and E. D. Cubuk, Scaling deep learning for materials discovery, Nature 10.1038/s41586-023-06735-9 (2023b).
@@
- Batatia et al. [2022] I. Batatia, D. P. Kovacs, G. N. C. Simm, C. Ortner, and G. Csanyi, MACE: Higher order equivariant message passing neural networks for fast and accurate force fields, in Advances in Neural Information Processing Systems, edited by A. H. Oh, A. Agarwal, D. Belgrave, and K. Cho (2022).
@@
- Kovács et al. [2023] D. P. Kovács, J. H. Moore, N. J. Browning, I. Batatia, J. T. Horton, V. Kapil, I.-B. Magdău, D. J. Cole, and G. Csányi, Mace-off23: Transferable machine learning force fields for organic molecules, arXiv preprint arXiv:2312.15211 (2023).

### v2

@@ Acknowledgements
We thank Nobuki Inoue and Tuan Minh Do for fruitful discussions. This project was supported by funding from the MEXT Quantum Leap Flagship Program (MEXTQLEAP) through Grant No. JPMXS0120319794, and the JST COI-NEXT Program through Grant No. JPMJPF2014. The completion of this research was partially facilitated by the JSPS Grants-in-Aid for Scientific Research (KAKENHI), specifically Grant Nos. JP23H03819 and JP21K18933. We thank the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo for the use of the facilities. This work was also achieved through the use of SQUID at the Cybermedia Center, Osaka University.
Notes and references
- Reymond et al. 2010 J.-L. Reymond, R. van Deursen, L. C. Blum and L. Ruddigkeit, MedChemComm, 2010, 1, 30.
@@
- Ruddigkeit et al. 2012 L. Ruddigkeit, R. van Deursen, L. C. Blum and J.-L. Reymond, Journal of Chemical Information and Modeling, 2012, 52, 2864–2875.
@@
- Gómez-Bombarelli et al. 2016 R. Gómez-Bombarelli, J. Aguilera-Iparraguirre, T. D. Hirzel, D. Duvenaud, D. Maclaurin, M. A. Blood-Forsythe, H. S. Chae, M. Einzinger, D.-G. Ha, T. Wu, G. Markopoulos, S. Jeon, H. Kang, H. Miyazaki, M. Numata, S. Kim, W. Huang, S. I. Hong, M. Baldo, R. P. Adams and A. Aspuru-Guzik, Nature Materials, 2016, 15, 1120–1127.
@@
- Curtarolo et al. 2013 S. Curtarolo, G. L. W. Hart, M. B. Nardelli, N. Mingo, S. Sanvito and O. Levy, Nature Materials, 2013, 12, 191–201.
- Nie et al. 2023 W. Nie, Q. Wan, J. Sun, M. Chen, M. Gao and S. Chen, Nature Communications, 2023, 14, 6671.
- Oganov et al. 2019 A. R. Oganov, C. J. Pickard, Q. Zhu and R. J. Needs, Nature Reviews Materials, 2019, 4, 331–348.
- Pierens 2014 G. K. Pierens, Journal of computational chemistry, 2014, 35, 1388–1394.
@@
- Hartman et al. 2016 J. D. Hartman, R. A. Kudla, G. M. Day, L. J. Mueller and G. J. O. Beran, Physical Chemistry Chemical Physics, 2016, 18, 21686–21709.
- Büning and Grimme 2023 J. B. K. Büning and S. Grimme, Journal of Chemical Theory and Computation, 2023, 19, 3601–3615.
- Chen and Ong 2022 C. Chen and S. P. Ong, Nature Computational Science, 2022, 2, 718–728.
- Lodewyk et al. 2011 M. W. Lodewyk, M. R. Siebert and D. J. Tantillo, Chemical Reviews, 2011, 112, 1839–1862.
- Lauro et al. 2020 G. Lauro, P. Das, R. Riccio, D. S. Reddy and G. Bifulco, The Journal of Organic Chemistry, 2020, 85, 3297–3306.
@@
- Hansen et al. 2015 K. Hansen, F. Biegler, R. Ramakrishnan, W. Pronobis, O. A. Von Lilienfeld, K.-R. Muller and A. Tkatchenko, The journal of physical chemistry letters, 2015, 6, 2326–2331.
- Deringer et al. 2021 V. L. Deringer, A. P. Bartók, N. Bernstein, D. M. Wilkins, M. Ceriotti and G. Csányi, Chemical Reviews, 2021, 121, 10073–10141.
@@
- Faber et al. 2017 F. A. Faber, L. Hutchison, B. Huang, J. Gilmer, S. S. Schoenholz, G. E. Dahl, O. Vinyals, S. Kearnes, P. F. Riley and O. A. von Lilienfeld, Journal of Chemical Theory and Computation, 2017, 13, 5255–5264.
@@
- Sajjan et al. 2022 M. Sajjan, J. Li, R. Selvarajan, S. H. Sureshbabu, S. S. Kale, R. Gupta, V. Singh and S. Kais, Chemical Society Reviews, 2022, 51, 6475–6573.
@@
- Keith et al. 2021 J. A. Keith, V. Vassilev-Galindo, B. Cheng, S. Chmiela, M. Gastegger, K.-R. Müller and A. Tkatchenko, Chemical Reviews, 2021, 121, 9816–9872.
- Wan et al. 2023 K. Wan, J. He and X. Shi, Advanced Materials, 2023, 2305758.
@@
- Wu et al. 2018 Z. Wu, B. Ramsundar, E. N. Feinberg, J. Gomes, C. Geniesse, A. S. Pappu, K. Leswing and V. Pande, Chemical Science, 2018, 9, 513–530.
@@
- Reiser et al. 2022 P. Reiser, M. Neubert, A. Eberhard, L. Torresi, C. Zhou, C. Shao, H. Metni, C. van Hoesel, H. Schopmans, T. Sommer et al., Communications Materials, 2022, 3, 93.
- Bartók et al. 2017 A. P. Bartók, S. De, C. Poelking, N. Bernstein, J. R. Kermode, G. Csányi and M. Ceriotti, Science advances, 2017, 3, e1701816.
- Kocer et al. 2022 E. Kocer, T. W. Ko and J. Behler, Annual Review of Physical Chemistry, 2022, 73, 163–186.
- Langer et al. 2022 M. F. Langer, A. Goeßmann and M. Rupp, npj Computational Materials, 2022, 8, 41.
- Liu et al. 2021 Z. Liu, L. Lin, Q. Jia, Z. Cheng, Y. Jiang, Y. Guo and J. Ma, Journal of Chemical Information and Modeling, 2021, 61, 1066–1082.
- Merchant et al. 2023 A. Merchant, S. Batzner, S. S. Schoenholz, M. Aykol, G. Cheon and E. D. Cubuk, Nature, 2023, 624, 80–85.
- Gao et al. 2020 P. Gao, J. Zhang, Q. Peng, J. Zhang and V.-A. Glezakou, Journal of Chemical Information and Modeling, 2020, 60, 3746–3754.
- Macalino et al. 2015 S. J. Y. Macalino, V. Gosu, S. Hong and S. Choi, Archives of Pharmacal Research, 2015, 38, 1686–1701.
@@
- Himanen et al. 2020 L. Himanen, M. O. Jäger, E. V. Morooka, F. Federici Canova, Y. S. Ranawat, D. Z. Gao, P. Rinke and A. S. Foster, Computer Physics Communications, 2020, 247, 106949.
- Gupta et al. 2021 A. Gupta, S. Chakraborty and R. Ramakrishnan, Machine Learning: Science and Technology, 2021, 2, 035010.
@@
- Gerrard et al. 2020 W. Gerrard, L. A. Bratholm, M. J. Packer, A. J. Mulholland, D. R. Glowacki and C. P. Butts, Chemical Science, 2020, 11, 508–515.
- Willatt et al. 2019 M. J. Willatt, F. Musil and M. Ceriotti, The Journal of Chemical Physics, 2019, 150, 154110.
- Bartók et al. 2013 A. P. Bartók, R. Kondor and G. Csányi, Physical Review B, 2013, 87, 184115.
- Faber et al. 2018 F. A. Faber, A. S. Christensen, B. Huang and O. A. von Lilienfeld, The Journal of Chemical Physics, 2018, 148, 241717.
@@
- Christensen et al. 2020 A. S. Christensen, L. A. Bratholm, F. A. Faber and O. Anatole von Lilienfeld, The Journal of Chemical Physics, 2020, 152, 044107.
@@
- Kabylda et al. 2023 A. Kabylda, V. Vassilev-Galindo, S. Chmiela, I. Poltavsky and A. Tkatchenko, nature communications, 2023, 14, 3562.
- Musil et al. 2021 F. Musil, A. Grisafi, A. P. Bartók, C. Ortner, G. Csányi and M. Ceriotti, Chemical Reviews, 2021, 121, 9759–9815.
- Bartók et al. 2010 A. P. Bartók, M. C. Payne, R. Kondor and G. Csányi, Physical review letters, 2010, 104, 136403.
@@
- Parsaeifard et al. 2021 B. Parsaeifard, D. Sankar De, A. S. Christensen, F. A. Faber, E. Kocer, S. De, J. Behler, O. Anatole von Lilienfeld and S. Goedecker, Machine Learning: Science and Technology, 2021, 2, 015018.
- Khan et al. 2023 D. Khan, S. Heinen and O. A. von Lilienfeld, The Journal of Chemical Physics, 2023, 159, 034106.
- Rupp et al. 2015 M. Rupp, R. Ramakrishnan and O. A. von Lilienfeld, The Journal of Physical Chemistry Letters, 2015, 6, 3309–3313.
@@
- Cordova et al. 2022 M. Cordova, E. A. Engel, A. Stefaniuk, F. Paruzzo, A. Hofstetter, M. Ceriotti and L. Emsley, The Journal of Physical Chemistry C, 2022, 126, 16710–16720.
@@
- Kohlhoff et al. 2009 K. J. Kohlhoff, P. Robustelli, A. Cavalli, X. Salvatella and M. Vendruscolo, Journal of the American Chemical Society, 2009, 131, 13894–13895.
- Tsitsvero et al. 2023 M. Tsitsvero, J. Pirillo, Y. Hijikata and T. Komatsuzaki, The Journal of Chemical Physics, 2023, 158, 194108.
- Paruzzo et al. 2018 F. M. Paruzzo, A. Hofstetter, F. Musil, S. De, M. Ceriotti and L. Emsley, Nature communications, 2018, 9, 4501.
- Fung et al. 2021 V. Fung, J. Zhang, E. Juarez and B. G. Sumpter, npj Computational Materials, 2021, 7, 84.
- Guan et al. 2021 Y. Guan, S. V. S. Sowndarya, L. C. Gallegos, P. C. S. John and R. S. Paton, Chemical Science, 2021, 12, 12012–12026.
@@
- Liu et al. 2019 S. Liu, J. Li, K. C. Bennett, B. Ganoe, T. Stauch, M. Head-Gordon, A. Hexemer, D. Ushizima and T. Head-Gordon, The Journal of Physical Chemistry Letters, 2019, 10, 4558–4565.
- Han and Choi 2021 H. Han and S. Choi, The Journal of Physical Chemistry Letters, 2021, 12, 3662–3668.
- Kwon et al. 2020 Y. Kwon, D. Lee, Y.-S. Choi, M. Kang and S. Kang, Journal of Chemical Information and Modeling, 2020, 60, 2024–2030.
- Jonas and Kuhn 2019 E. Jonas and S. Kuhn, Journal of cheminformatics, 2019, 11, 1–7.
@@
- Schütt et al. 2017 K. T. Schütt, P.-J. Kindermans, H. E. Sauceda, S. Chmiela, A. Tkatchenko and K.-R. Müller, SchNet: A continuous-filter convolutional neural network for modeling quantum interactions, 2017.
- Chen et al. 2019 C. Chen, W. Ye, Y. Zuo, C. Zheng and S. P. Ong, Chemistry of Materials, 2019, 31, 3564–3572.
- Gasteiger et al. 2021 J. Gasteiger, F. Becker and S. Günnemann, Advances in Neural Information Processing Systems, 2021, 34, 6790–6802.
- Liao and Smidt 2023 Y.-L. Liao and T. Smidt, Equiformer: Equivariant Graph Attention Transformer for 3D Atomistic Graphs, 2023.
- Liao et al. 2024 Y.-L. Liao, B. Wood, A. Das and T. Smidt, EquiformerV2: Improved Equivariant Transformer for Scaling to Higher-Degree Representations, 2024.
- Xu et al. 2019 K. Xu, W. Hu, J. Leskovec and S. Jegelka, How Powerful are Graph Neural Networks?, 2019.
@@
- Stärk et al. 2022 H. Stärk, D. Beaini, G. Corso, P. Tossou, C. Dallago, S. Günnemann and P. Lió, Proceedings of the 39th International Conference on Machine Learning, 2022, pp. 20479–20502.
- Jiang et al. 2021 D. Jiang, Z. Wu, C.-Y. Hsieh, G. Chen, B. Liao, Z. Wang, C. Shen, D. Cao, J. Wu and T. Hou, Journal of cheminformatics, 2021, 13, 1–23.
- Kang et al. 2020 S. Kang, Y. Kwon, D. Lee and Y.-S. Choi, Journal of Chemical Information and Modeling, 2020, 60, 3765–3769.
- Takamoto et al. 2022 S. Takamoto, C. Shinagawa, D. Motoki, K. Nakago, W. Li, I. Kurata, T. Watanabe, Y. Yayama, H. Iriguchi, Y. Asano et al., Nature Communications, 2022, 13, 2991.
@@
- Musaelian et al. 2023 A. Musaelian, S. Batzner, A. Johansson, L. Sun, C. J. Owen, M. Kornbluth and B. Kozinsky, Nature Communications, 2023, 14, 579.
@@
- Batatia et al. 2022 I. Batatia, D. P. Kovacs, G. N. C. Simm, C. Ortner and G. Csanyi, Advances in Neural Information Processing Systems, 2022.
@@
- Kovács et al. 2023 D. P. Kovács, J. H. Moore, N. J. Browning, I. Batatia, J. T. Horton, V. Kapil, I.-B. Magdău, D. J. Cole and G. Csányi, arXiv preprint arXiv:2312.15211, 2023.